# Memory preservation and cooperative shielding in complex quantum networks

### v2

@@ 3.4 Dependence on initial conditions
We start by seeking an understanding of the importance of initial conditions, as we have done in Section 2. Since the analytical machinery that we established for regular networks seems to not apply for irregular graphs, we seek for alternatives. It turns out that this is possible at least for the short-time IPR behavior, by means of the so-called cascade model introduced in [79]: in essence, this monitors the initial state $|s(0)\rangle$ coupling with other basis elements at different moments $|s(t)\rangle$.
@@
The underlying idea is as follows, see Fig. 5(a)(a) for the concept. At $t=0$, only the subset of state vectors $\mathcal{F}_{0}$ containing $|s\rangle$ participates in the temporal evolution. Immediately after, to first-order approximation the temporal evolution operator connects the initial state to the basis vectors $|f_{i}\rangle$ for which the matrix elements $\langle f_{i}|\hat{\mathcal{H}}|s\rangle$ are non-vanishing, thus extending the subset of involved states to $\mathcal{F}_{1}$. More precisely, the subset of states grows as
$\mathcal{F}_{0}=\{|s\rangle\}\xrightarrow{\exp{(\mathrm{i}\hat{\mathcal{H}}t)}\simeq(\hat{\mathbb{I}}-\mathrm{i}\hat{\mathcal{H}}t)}\mathcal{F}_{1}=\{|s\rangle\}\cup\{|f_{i}\rangle\,,\,\text{$f_{i}$ first neighbors of $s$}\}\,.$ (12)
@@
In fact, one can set and resolve rate equations for the probabilities of finding the state evolution $|s(t)\rangle$ to a particular basis element $|f_{i}\rangle$:
$P_{f_{i}}(t)=|\langle f_{i}|s(t)\rangle|^{2}\,.$
Borgonovi et al. have found a transparent expression for the evolution of IPR at short times, in the approximation where any back-flow to the initial site can be neglected [79]. This is:
@@
where the sum is performed over the first-neighbor sites of the initial one, and $\mathcal{N}_{1}$ is the number of such nodes plus one (referring to vertex $s$). Let us better understand Eq. 13: as the survival probability $P_{s}$ decreases because of the excitation spreading, the initial site couples with its first neighbors in the interaction topology. In other words, the IPR evolution is strictly dependent on the graph upon which it lives, through its adjacency matrix $A_{ij}$ describing the existing connections and being proportional to the XX Hamiltonian itself, when written in the restricted basis of computational vectors having one excitation. Notice that $A_{ij}=1$ if sites $i$ and $j$ are linked, else $A_{ij}=0$. Under a different perspective, the degree of the initial node (i.e. the number of first neighbors in the graph) governs the transport onset at short times.
@@
Another illuminating way of looking at the impact of initial conditions on excitation spreading regards increasing connectivities close to the limit of all-to-all connected networks. The results are illustrated in Fig. 6(ab)(a-b), where we consider a network with a given 2k=L22k=L-2 parameter, closest to all-to-all case but not-fully connected, and impose a non-zero rewiring probability providing the sites with different degrees, i.e. different number of links with other elements. Then, we compute the IPR by picking different initial sites according to their degree. We do this for both long, Fig. 6 $(a)$, and short, Fig. 6 $(b)$ temporal windows, and we observe that the IPR time-evolution behavior for the all-to-all case soon breaks into the general small-world system behavior. Here, the degeneracy lifting due to the symmetry reduction with respect to the all-to-all case makes the IPR decreasing over time. The stationary value of the IPR critically depends on the degree of connectivity of the initial site.
@@
Figure 6: Understanding transport in highly-connected networks. $(a)$ IPR behaviors for long periods of time for varying initial sites, denoted by their degree, in a fixed network with $L=33$, 2k=L22k=L-2, $\beta=0.1$. $(b)$ Same as $(a)$ but for short times; here $\omega^{*}=JL/2=33$ dictates the oscillation periods of the all-to-all connected IPR. $(c)$ IPR distance (14) between all-to-all and small-world networks graphs as a function of the initial site’s degree related to L1L-1 at increasing times as in the legend.
3.5 Infinite-time localization depending on graph properties
@@
We summarize the results in Fig. 7. The boxplot, Fig. 7(a)(a), displays the average IPR as a function of the number of initial neighbors $2k$, for different values of the rewiring probability $\beta$. We see that localization, i.e. higher IPR values, is favored for extremal values of $k$ (1 and 15). The lower extreme corresponds to the obvious case of low connectivity, specifically similar to a nearest-neighbor connection. The latter, instead, corresponds to the case of highly-connected graphs with significant interference effects. Delocalization manifests instead at intermediate connectivity conditions. The rewiring probability $\beta$ does not significantly influence the average value, essentially dominated by the number of neighbors. We associate variations in this behavior with the limited number of sampled graphs.
@@
We now represent in Fig. 7(b)(b) the same IPR results in a scatter plot where the graphs are characterized by their average distance and average clustering instead, see Appendix C. We see that relatively localized long-time states are possible away from the limit of all-to-all connected graphs, that is with average clustering and average distance equal to 1. Quite interestingly, non-trivial steady states emerge not only in the highly connected cases but also in sparsely connected graphs with high-average distance (lower-right part of the plot), due to the slow propagation of excitation and small interference. This behavior heralds the potential for a memory-preserving property of the initial quench even away from the all-to-all case, should a certain decoding protocol be provided. In contrast, with $k$ values comparable to the all-to-all case with L(L1)/2L(L-1)/2 (upper-left part), signatures of cooperative shielding appear, typical of complete connection. The difference between the finite IPR values (green to light-blue symbols) shown in the top-left and in the bottom-right part of Fig. 7(b)(b) is better envisioned by inspecting Fig. 7(c)(c). Here we show a typical IPR vs. time $tJ/2$ behavior (blue line) for parameter values in the bottom-right part of $(b)$ for $k=1$. The IPR signal oscillates around an average (dot-dashed horizonatla line) that is clearly above the threshold $1/L$ (dashed horizontal line) of complete delocalization, though the IPR has a lower value with respect to the IPR in a corresponding all-to-all case (yellow). The cascade model prediction (red) only provides a reasonable fit at all times. The differences are due to the fact that the cascade model neglects all backflow processes. Thus, we might infer that the emerging localization in this region is enhanced by quantum interference effects among different paths, in a mechanism similar to Anderson localization [80]. Finally we note that the question of whether cooperative shielding survives in more general networks deserves further study which requires a scaling analysis as a function of the system size. This topic will be analyzed in a future work.
@@
Figure 7: Effect of network topology on the long-time average IPR. For each pair $(k,\beta)$ of initial conditions in the graph generation, the computed average IPR in the time window $tJ/2\in[90,100]$, to ensure that the excitation could travel across the whole graph. Each pair is tested 20 times and each time the initial site is chosen at random. $(a)$ Long-time averaged IPR vs. the number of initial neighbors $2k$ normalized to L1L-1. The horizontal dashed line on top is the average IPR for the all-to-all case. For each 2k/(L1)2k/(L-1) value, boxes represent the data’s interquartile range, i.e. all points whose mean IPR belong between the 25th and 75th quartile, and have a horizontal line marking the median; whisker lines extend up to 1.5 times the interquartile range outside of boxes; crosses depict the outliers, or points that are not included in neither boxes or whiskers; finally, the red diamonds display the average IPR sampling. Localization is seen to occur at extremal values of $k$, i.e. nearest-neighbor and all-to-all type of connectivity. Parameter values: $k=1,\dots,15$, $\beta=0.1,0.2,\dots,1$, $L=33$. $(b)$ Long-time averaged IPR in the form of a scatter plot in the parameter space of average clustering $\mathcal{C}$ and average distance $\mathcal{D}$ (data having $\mathcal{C}=0$ is not shown). The general trend going from lower-right to upper-left denote increasing values of $k$. Notice that, away from all-to-all connectivity with $\mathcal{C}=\mathcal{D}=1$, some degree of localization is still possible for more regular networks, i.e. sparse graphs with large $\mathcal{D}$ and intermediate to very low values of clustering $\mathcal{C}$. We see this in panel $(c)$ showing a typical graph of IPR vs. dimensionless time $tJ/2$ (blue line) for parameter values $k=1,\beta=0.5,\mathcal{D}=7.0,\mathcal{C}=0$ in the bottom-right part of $(b)$, corresponding to the network in $(d)$ initialized at site $17$, as depicted. The IPR displays irregular oscillations around an average (dashed line) that is well above the threshold $1/L$ of complete delocalization. For comparison, the IPR signal is shown with the respective cascade model (red).
3.6 Power-law-dependent hopping
@@
A different way to break space regularity in networks is to consider exchange coupling that vary with physical distance, specifically characterized by a power law, i.e. $J_{ij}\propto\text{dist}(i,j)^{-\alpha}$ in Eq. (1) which in the case of PBC is equivalent to min(|ij|,L|ij|)(|i-j|,L-|i-j|). This test case constitutes a good bench to showcase cooperative shielding, introduced in Section 1, when considering different initial conditions: in fact, on the one hand increasing values of the power law coefficient cause shorter ranges for the interactions (as was for diverse combinations of $k$ and $\beta$ for small-worlds), and on the other hand each site is equivalent to the others (unlike the previous varying network case). Here we analyze the memory effects in such models and the dynamics focusing on the cooperative shielding effects.
3.6.1 Memory effects
@@
Here we analyze the memory-preserving effect discussed in Section 2.1 in networks with power-law interactions, starting with a single excitation and taking the Fourier transform of $S^{z}_{\text{initial}}(t)$, see Fig. 8(a)(a). As we increase $\alpha$ the high-energy frequencies are reduced but they still persist. Also in this case, this result could allow us to extend the notion of memory-preserving effect to generic range of hopping.
3.6.2 Dynamics/shielding
@@
The results are displayed in Fig. (9)$(b)$, where we obtain $\Delta\tau_{1/2}$ for different values of $\alpha$ and different scales or number of sites $L$. We first notice that the stopping time difference is a decreasing function of $\alpha$ and reaches the lower bound corresponding to the nearest-neighbor network case already for $\alpha\gtrsim 5$.
@@
The results shown in this section are consistent with what was claimed in [71] where it was shown that single excitation transport in presence of long range hopping is highly dependent on the initial conditions. The results presented here confirm that this is a general effect of long-range hopping models. Indeed for short-range hopping, $\alpha\gg 1$ no dependence on the initial condition is shown, see Fig. 9(a)(a). Note that the antisymmetric initial state has a large overlap with the shielded subspace [71], even if it does not lie completely in the shielding subspace and this could explain the lack of cooperativity (dependence on the size of the system) shown in Fig. 9(b)(b). The data has been detrended before the Fourier transform by removing the zero-frequency component.
@@
4 Discussion and conclusions
This work is motivated by the relevance of complex quantum networks in the modeling of biological systems and as a powerful tool to extract relevant information for many-body systems. We have therefore addressed the transport properties of quantum networks as governed by different parameters such as the connectivity graph, the topology and the range of the couplings connecting the nodes. These parameters can be tuned in current quantum technology platforms [14]. We have focused on networks described by the XXZ Hamiltonian, that is a paradigmatic model with a variety of applications [74]. To gain insights on this complex matter, we have adopted a methodology that progressively develops understanding from the simpler to harder cases: we start from the all-to-all connectivity with constant coupling among the nodes, then introducing power-law interactions in the regular network, and eventually passing to irregular, complex networks. The methodology employs combined analytical and numerical methods, revolving around the cooperative shielding concept and quantum cascade model [70] accompanied by exact diagonalization. The main focus of this manuscript is on the effect of long range interactions and connectivity on the dynamics of the network. We show that long range coupling in highly connected networks imposes strong constraints on the dynamics which manifest themselves in the memory-preserving and cooperative shielding effect [70, 71].
Motivated by the findings in [5], we have analyzed the robust preservation of the memory of the number of initial excitations over long times. This peculiar phenomenon essentially results into a spatiotopic representation of the number of frequencies in the power spectrum of the network, that are as many as the number of excitations initially injected. Even more interestingly, the uncertainty in the number of excitations that the quantum network can map, linearly increases with the number itself. This behavior is shown in [5] to elegantly account for the so-called Weber’s law for numerosity perception in neuroscience problems, after mapping the essential functions of the classical neuronal network into an open quantum network (resulting into the XXZ model), and in contrast with the poor performance of conventional biologically inspired models like artificial neural networks. In fact, this memory preserving property can be of much interest for concrete implementations with current quantum technologies.

### v3

@@ III.4 Dependence on initial conditions
We start by seeking an understanding of the importance of initial conditions, as we have done in Section II. Since the analytical machinery that we established for regular networks seems to not apply for irregular graphs, we seek for alternatives. It turns out that this is possible at least for the short-time IPR behavior, by means of the so-called cascade model introduced in [borgonovi2019ipr]: in essence, this monitors the initial state $|x(0)\rangle=|s\rangle$ coupling with other basis elements at different moments $|x(t)\rangle$; here, $|s\rangle$ denotes the initial site of the excitation.
@@
The underlying idea is as follows, see Fig. 5(a)(a) for the concept. At $t=0$, only the subset of state vectors $\mathcal{F}_{0}$ containing $|s\rangle$ participates in the temporal evolution. Immediately after, to first-order approximation the temporal evolution operator connects the initial state to the basis vectors $|f_{i}\rangle$ for which the matrix elements $\langle f_{i}|\hat{\mathcal{H}}|s\rangle$ are non-vanishing, thus extending the subset of involved states to $\mathcal{F}_{1}$. More precisely, the subset of states grows as:
$\mathcal{F}_{0}$ $=\{|s\rangle\}\xrightarrow{\exp{(\mathrm{i}\hat{\mathcal{H}}t)}\simeq(\hat{\mathbb{I}}-\mathrm{i}\hat{\mathcal{H}}t)}\mathcal{F}_{1}$ (13)
$=\{|s\rangle\}\cup\{|f_{i}\rangle\,,\,\text{$f_{i}$ first neighbors of $s$}\}\,.$
@@
In fact, one can set and resolve rate equations for the probabilities of finding the state evolution $|x(t)\rangle$ to a particular basis element $|f_{i}\rangle$:
$P_{s}(t)$ $=|\langle s|x(t)\rangle|^{2}\,$,
$P_{f_{i}}(t)$ $=|\langle f_{i}|x(t)\rangle|^{2}\,.$
Borgonovi et al. have found a transparent expression for the evolution of IPR at short times, in the approximation where any back-flow to the initial site can be neglected [borgonovi2019ipr]. This is:
@@
where the sum is performed over the first-neighbor sites of the initial one, and $\mathcal{N}_{1}$ is the number of such nodes plus one (referring to vertex $s$). Let us better understand Eq.(14): as the survival probability $P_{s}$ decreases because of the excitation spreading, the initial site couples with its first neighbors in the interaction topology. In other words, the IPR evolution is strictly dependent on the graph upon which it lives, through its adjacency matrix $A_{ij}$ describing the existing connections and being proportional to the XX Hamiltonian itself, when written in the restricted basis of computational vectors having one excitation. Notice that $A_{ij}=1$ if sites $i$ and $j$ are linked, else $A_{ij}=0$. Under a different perspective, the degree of the initial node (i.e. the number of first neighbors in the graph) governs the transport onset at short times.
@@
Another illuminating way of looking at the impact of initial conditions on excitation spreading regards increasing connectivities close to the limit of all-to-all connected networks. The results are illustrated in Fig. 6(ab)(a-b), where we consider a network with a given $\bar{g}=L-3$ parameter, closest to all-to-all case but not-fully connected, and impose a non-zero rewiring probability providing the sites with different degrees, i.e. different number of links with other elements. Then, we compute the IPR by picking different initial sites according to their degree. We do this for both long, Fig. 6 $(a)$, and short, Fig. 6 $(b)$ temporal windows, and we observe that the IPR time-evolution behavior for the all-to-all case soon breaks into the general small-world system behavior. Here, the degeneracy lifting due to the symmetry reduction with respect to the all-to-all case makes the IPR decreasing over time. The stationary value of the IPR critically depends on the degree of connectivity of the initial site.
@@
Figure 6: Understanding transport in highly-connected networks. $(a)$ IPR behaviors for long periods of time for varying initial sites, denoted by their degree, in a fixed network with $L=33$, $\bar{g}=L-3$, $\beta=0.1$. $(b)$ Same as $(a)$ but for short times; here $\omega^{*}=JL/2=33$ dictates the oscillation periods of the all-to-all connected IPR. $(c)$ IPR distance (15) between all-to-all and small-world networks graphs as a function of the initial site’s degree related to L1L-1 at increasing times as in the legend.
III.5 Infinite-time localization depending on graph properties
@@
We summarize the results in Fig. 7. The boxplot, Fig. 7(a)(a), displays the average IPR as a function of the number of initial neighbors $\bar{g}$, for different values of the rewiring probability $\beta$. We see that localization, i.e. higher IPR values, is favored for extremal values of $k$ (1 and 15). The lower extreme corresponds to the obvious case of low connectivity, specifically similar to a nearest-neighbor connection. The latter, instead, corresponds to the case of highly-connected graphs with significant interference effects. Delocalization manifests instead at intermediate connectivity conditions. The rewiring probability $\beta$ does not significantly influence the average value, essentially dominated by the number of neighbors. We associate variations in this behavior with the limited number of sampled graphs.
@@
We now represent in Fig. 7(b)(b) the same IPR results in a scatter plot where the graphs are characterized by their average distance and average clustering instead, see Appendix D. We see that relatively localized long-time states are possible away from the limit of all-to-all connected graphs, that is with average clustering and average distance equal to 1. Quite interestingly, non-trivial steady states emerge not only in the highly connected cases but also in sparsely connected graphs with high-average distance (lower-right part of the plot), due to the slow propagation of excitation and small interference. This behavior heralds the potential for a memory-preserving property of the initial quench even away from the all-to-all case, should a certain decoding protocol be provided. In contrast, with $k$ values comparable to the all-to-all case with L(L1)/2L(L-1)/2 (upper-left part), signatures of cooperative shielding appear, typical of complete connection. The difference between the finite IPR values (green to light-blue symbols) shown in the top-left and in the bottom-right part of Fig. 7(b)(b) is better envisioned by inspecting Fig. 7(c)(c). Here we show a typical IPR vs. time $tJ/2$ behavior (blue line) for parameter values in the bottom-right part of $(b)$ for $k=1$. The IPR signal oscillates around an average (dot-dashed horizontal line) that is clearly above the threshold $1/L$ (dashed horizontal line) of complete delocalization, though the IPR has a lower value with respect to the IPR in a corresponding all-to-all case (yellow). The cascade model prediction (red) only provides a reasonable fit at all times. The differences are due to the fact that the cascade model neglects all backflow processes. Thus, we might infer that the emerging localization in this region is enhanced by quantum interference effects among different paths, in a mechanism similar to Anderson localization [alet2018_mbl]. Finally we note that the question of whether cooperative shielding survives in more general networks deserves further study which requires a scaling analysis as a function of the system size. This topic will be analyzed in a future work.
@@
Figure 7: Effect of network topology on the long-time average IPR. For each pair $(k,\beta)$ of initial conditions in the graph generation, the computed average IPR in the time window $tJ/2\in[90,100]$, to ensure that the excitation could travel across the whole graph. Each pair is tested 20 times and each time the initial site is chosen at random. $(a)$ Long-time averaged IPR vs. the number of initial neighbors $\bar{g}$ normalized to L1L-1. The horizontal dashed line on top is the average IPR for the all-to-all case. For each $\bar{g}/(L-1)$ value, boxes represent the data’s interquartile range, i.e. all points whose mean IPR belong between the 25th and 75th quartile, and have a horizontal line marking the median; whisker lines extend up to 1.5 times the interquartile range outside of boxes; crosses depict the outliers, or points that are not included in neither boxes or whiskers; finally, the red diamonds display the average IPR sampling. Localization is seen to occur at extremal values of $k$, i.e. nearest-neighbor and all-to-all type of connectivity. Parameter values: $k=1,\dots,15$, $\beta=0.1,0.2,\dots,1$, $L=33$. $(b)$ Long-time averaged IPR in the form of a scatter plot in the parameter space of average clustering $\mathcal{C}$ and average distance $\mathcal{D}$ (data having $\mathcal{C}=0$ is not shown). The general trend going from lower-right to upper-left denote increasing values of $k$. Notice that, away from all-to-all connectivity with $\mathcal{C}=\mathcal{D}=1$, some degree of localization is still possible for more regular networks, i.e. sparse graphs with large $\mathcal{D}$ and intermediate to very low values of clustering $\mathcal{C}$. We see this in panel $(c)$ showing a typical graph of IPR vs. dimensionless time $tJ/2$ (blue line) for parameter values $k=1,\beta=0.5,\mathcal{D}=7.0,\mathcal{C}=0$ in the bottom-right part of $(b)$, corresponding to the network in $(d)$ initialized at site $17$, as depicted. The IPR displays irregular oscillations around an average (dashed line) that is well above the threshold $1/L$ of complete delocalization. For comparison, the IPR signal is shown with the respective cascade model (red).
III.6 Power-law-dependent hopping
@@
A different way to break space regularity in networks is to consider exchange coupling that vary with physical distance, specifically characterized by a power law, i.e. $J_{ij}\propto\text{dist}(i,j)^{-\alpha}$ in Eq.(1) which in the case of PBC is equivalent to min(|ij|,L|ij|)(|i-j|,L-|i-j|). This test case constitutes a good bench to showcase cooperative shielding, introduced in Section I, when considering different initial conditions: in fact, on the one hand increasing values of the power law coefficient cause shorter ranges for the interactions (as was for diverse combinations of $k$ and $\beta$ for small-worlds), and on the other hand each site is equivalent to the others (unlike the previous varying network case). Here we analyze the memory effects in such models and the dynamics focusing on the cooperative shielding effects.
III.6.1 Memory effects
@@
Here we analyze the memory-preserving effect discussed in Section II.1 in networks with power-law interactions, starting with a single excitation and taking the Fourier transform of $S^{z}_{\text{initial}}(t)$, see Fig. 8(a)(a). As we increase $\alpha$ the high-energy frequencies are reduced but they still persist. Also in this case, this result could allow us to extend the notion of memory-preserving effect to generic range of hopping.
III.6.2 Dynamics/shielding
@@
The results are displayed in Fig. 9(b)(b), where we obtain $\Delta\tau_{1/2}$ for different values of $\alpha$ and different scales or number of sites $L$. We first notice that the stopping time difference is a decreasing function of $\alpha$ and reaches the lower bound corresponding to the nearest-neighbor network case already for $\alpha\gtrsim 5$.
@@
The results shown in this Section are consistent with what was claimed in [celardo2016_shielding] where it was shown that single excitation transport in presence of long-range hopping is highly dependent on the initial conditions. The results presented here confirm that this is a general effect of long-range hopping models. Indeed for short-range hopping, $\alpha\gg 1$ no dependence on the initial condition is shown, see Fig. 9(a)(a). Note that the antisymmetric initial state has a large overlap with the shielded subspace [celardo2016_shielding], even if it does not lie completely in the shielding subspace and this could explain the lack of cooperativity (dependence on the size of the system) shown in Fig. 9(b)(b). The data has been detrended before the Fourier transform by removing the zero-frequency component.
@@
IV Discussion and conclusions
This work is motivated by the relevance of complex quantum networks in the modeling of biological systems and as a powerful tool to extract relevant information for many-body systems. We have therefore addressed the transport properties of quantum networks as governed by different parameters such as the connectivity graph, the topology and the range of the couplings connecting the nodes. These parameters can be tuned in current quantum technology platforms [yagomalo2024review]. We have focused on networks described by the XXZ Hamiltonian, that is a paradigmatic model with a variety of applications [franchini2017integrable]. To gain insights on this complex matter, we have adopted a methodology that progressively develops understanding from the simpler to harder cases: we start from the all-to-all connectivity with constant coupling among the nodes, then introducing power-law interactions in the regular network, and eventually passing to irregular, complex networks. The methodology employs combined analytical and numerical methods, revolving around the cooperative shielding concept and quantum cascade model [santos2016cooperative] accompanied by Exact Diagonalization. The main focus of this manuscript is on the effect of long-range interactions and connectivity on the dynamics of the network. We show that long-range coupling in highly connected networks imposes strong constraints on the dynamics which manifest themselves in the memory-preserving and cooperative shielding effect [santos2016cooperative, celardo2016_shielding].
Motivated by the findings in [YagoMalo2023], we have analyzed the robust preservation of the memory of the number of initial excitations over long times. This peculiar phenomenon essentially results into a spatiotopic representation of the number of frequencies in the power spectrum of the network, that are as many as the number of excitations initially injected. Even more interestingly, the uncertainty in the number of excitations that the quantum network can map, linearly increases with the number itself. This behavior is shown in [YagoMalo2023] to elegantly account for the so-called Weber’s law for numerosity perception in neuroscience problems, after mapping the essential functions of the classical neuronal network into an open quantum network (resulting into the XXZ model), and in contrast with the poor performance of conventional biologically inspired models like artificial neural networks. In fact, this memory preserving property can be of much interest for concrete implementations with current quantum technologies.